A SOCCER PLAYER GIVES A 2KG BALL A HORIZONTAL VELOCITY OF 20M/S. THE BALL SLIDES ALONG THE TURF AGAINST A FRICTIONAL FORCE OF 4N,WHAT IS THE VELOCITY OF THE BALL AFTER 25 M?


1
Expert's answer
2021-03-23T08:14:12-0400

According to the work-energy theorem, the work done by the frictional force is equal to the negative ball's kinetic energy change. By definition, the work is:


"W = Fs"

where "F = 4N" is the frictional force, and "s = 25m" is the length of the ball's path.

The initial kinetic energy is:


"K_i = \dfrac{mv_i^2}{2}"

where "m = 2kg" is the mass of the ball, and "v_i = 20m\/s" is its initial speed. Thus, obtain:


"K_i - K_f = W\\
K_f = K_i -W = \dfrac{mv_i^2}{2}-Fs"

On the other hand, the final kinetic energy "K_f" is:


"K_f = \dfrac{mv_f^2}{2}"

where "v_f" is the final speed of the ball. Thus, obtain:


"\dfrac{mv_f^2}{2}=\dfrac{mv_i^2}{2}-Fs\\
v_f^2 = v_i^2-\dfrac{2Fs}{m}\\
v_f = \sqrt{ v_i^2-\dfrac{2Fs}{m}}\\
v_f = \sqrt{ 20^2-\dfrac{2\cdot 4\cdot 25}{2}} \approx 17.3 m/s"

Answer. 17.3 m/s.